If f(x)=cos(x)-sin(x) then what is the value of f(1)?

Grade:11

2 Answers

Arun
6 years ago
here in f(1), 1 is angle in radian.
so f(1)= cos1−sin1
this will be a negative answer as in [0,\pi/4) cos will be greater than sin.
and 1 will be less than \pi/4.
Vikas TU
6 years ago
Dear Student,
f(x) = cosx – sinx.
f(1) = cos1 – sin1 = 0.54 – 0.84 = -0.3
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)